If the 4th term of an arithmetic sequence is 3 and the 20th term is 35, find the 100th term of the sequence.
I believe the equation = -5 + 2n
and maybe the 100th number is 95
It does not make senses to me.

Answer by rothauserc(4718)

An arithmetic sequence uses the following formula for the nth tern
Xn = a +(d*(n-1)), where Xn is the nth term, a is the first term, n is number of the term in the sequence, we are given
3 = a +(d*(4-1)) = a + 3d
35 = a +(d*(20-1)) = a +19d
using the first equation solve for a
a = 3-3d
substitute for a in second equation
35 = 3-3d +19d
16d = 32
d = 2 and a = 3 - 6 = -3
******************************
our equation for the nth term is
Xn = -3 +(2*(n-1))
******************************
now for the 100th term
X100 = -3 +(2*(100-1)) = 195